OpenStudy (anonymous):

Which equation is of a circle that has a center at (3, -2) and a radius of 9?

OpenStudy (anonymous):

(x + 3)2 + (y - 2)2 = 9 (x + 3)2 + (y - 2)2 = 81 (x - 3)2 + (y + 2)2 = 9 (x - 3)2 + (y + 2)2 = 81

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ (x-h)^2+(y-k)^2 +r^2 }\) where \(\large\color{black}{ (h,k) }\) is the center, and \(\large\color{black}{ r }\) is the radius.

OpenStudy (solomonzelman):

I meant, \(\large\color{black}{ (x-h)^2+(y-k)^2 =r^2 }\) and so, \(\large\color{black}{ (x-(3)~)^2+(y-(-2)~)^2 =9^2 }\)

OpenStudy (solomonzelman):

can you finish it?

OpenStudy (anonymous):

How do i finish it ?

OpenStudy (solomonzelman):

can you finish it?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

But i thinking it is the last answer choice ?

OpenStudy (solomonzelman):

\(\large\color{black}{ a-(-b)=a+b }\)

OpenStudy (solomonzelman):

yes, it is \(\large\color{black}{ (x-3)~)^2+(y+2)^2 =81 }\)

OpenStudy (anonymous):

Thank you :) i am going to graduate

OpenStudy (solomonzelman):

Lol, :) yw

OpenStudy (anonymous):

I will let you know if i have other questions sir :) ;)